OpenStudy (anonymous):

Find all solutions to the equation. 2sin theta + 1 = 0 Write your answer in radians in terms of pie , and use "or" as necessary.

OpenStudy (anonymous):

isolate theta

OpenStudy (anonymous):

\(\huge \theta\)

OpenStudy (anonymous):

start with \[\sin(\theta)=-\frac{1}{2}\] then figure out what numbers have \(-\frac{1}{2}\) as their sine

OpenStudy (anonymous):

Ok thanks one other thing though. Im not sure how to put my answer in this format. This is the example answer they gave. Example: /Users/Amir/Desktop/NM.gif

OpenStudy (anonymous):

they want you to format the answer like this. Example θ = (⅔π + nπ) radians and θ = (5π/3 + nπ) radians . . . where n is an integer.

OpenStudy (anonymous):

Im not sure what to put in for the "n" value